OpenStudy (anonymous):

Guys how would i prove this (cot(x-1))/(cot(x+1)) = (cos2x)/(1+sin2x)

OpenStudy (anonymous):

Theta and x are same

OpenStudy (anonymous):

can u coorect ur question ...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

cos2x= (cosx)^2 - (sinx)^2 =(cosx+sinx)(cosx-sinx) and 1+sin2x= (cosx)^2 + (sinx)^2 +2sinxcosx=(cosx+sinx)^2 hence RHS = (cos2x)/(1+sin2x) =((cosx)^2 - (sinx)^2 ) / ( (cosx)^2 + (sinx)^2 +2sinxcosx ) = (cosx+sinx)(cosx-sinx)/(cosx+sinx)^2 = (cosx-sinx)/(cosx+sinx) =((cosx-sinx)/sinx) / ((cosx+sinx)/sinx) [dividing both num and denom by sinx) =(cosx/sinx - sinx/sinx) / (cosx/sinx + sinx/sinx) =(cotx -1)/(cotx+1)

OpenStudy (anonymous):

so ur LHS hopefully is =(cotx -1)/(cotx+1) to prove form LHS to RHS ujust rewrite the above in reverse way

OpenStudy (anonymous):

Oh wow thanks ya i miswrote the original problem. If you don't mind me asking, how exactly did u see that i had miswrote it?

OpenStudy (anonymous):

u r learning compound angles i suppose..

OpenStudy (anonymous):

haha yea. What exactly is the first thing you did to the bottom?

OpenStudy (anonymous):

nvm i see what you did

OpenStudy (anonymous):

didn't get u..

OpenStudy (anonymous):

Hey Thank You so much

OpenStudy (anonymous):

welcome